OpenStudy (anonymous):

evaluate 10!/6! A) 5040 B) 604800 C) 720 D) 3628800

jhonyy9 (jhonyy9):

10! 6!*7*8*9*10 ---- = ----------- = 7*8*9*10 = 56*90=5040 6! 6! this mean that A). is right sure